OpenStudy (anonymous):

find the slope intercept form for the line satifying the following conditions x intercept 2, y intercept 1/5

OpenStudy (lgbasallote):

hint: use the intercept form \[\frac xa + \frac yb = 1\] where: a = x-intercept b = y-intercept does that help?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

too bad. would have been useful

OpenStudy (anonymous):

How should I set up the equation for this problem ?

OpenStudy (anonymous):

That method above will work great.

OpenStudy (anonymous):

\[\frac{x}{2}+\frac{y}{\frac{1}{5}}=1\]

OpenStudy (anonymous):

Or... if you use point slope equation with the two points you were given (2, 0) and (0, 1/5) and the slope equation to get m.... but that will take a lot more steps.

OpenStudy (anonymous):

I would just take that first equation and manipulate it to be y=mx+b. As long as it is simplified, it will take the right form as y=(everything else)